# Sample Complexity Analysis of Multi-Target Detection via Markovian and Hard-Core Multi-Reference Alignment

### v2

@@ Appendix A Proofs
For the i.i.d. measurements ${Y}^{(M)}$, we have the standard bias-variance decomposition
@@
Now, for the dependent data $Z^{(M)}$, we can estimate the bias of the estimator $\widehat{F}(Z^{(M)})$ as
@@
$\leq\left|\dfrac{1}{|\mathcal{I}_{M}|^{2}}\sum_{i\in\mathcal{I}_{M}}\operatorname{Var}\left(F(Z_{i})\right)-\operatorname{Var}\left(\widehat{F}(Y^{(M)})\right)\right|$
(61) $\quad+\dfrac{1}{|\mathcal{I}_{M}|^{2}}\sum_{\begin{subarray}{c}(i,j)\in\mathcal{I}_{M}^{2}\\ i\neq j\end{subarray}}\left|\operatorname{Cov}\left(F(Z_{i}),F(Z_{j})\right)\right|.$
Since ${Y}^{(M)}=\{{Y}_{i}\}_{1\leq i\leq M^{d}}$ are i.i.d., we have
@@
Next, we prove that, for any $(i,j)\in\mathcal{I}_{M}^{2}$ with $i\neq j$, it holds that
@@
Then, combining (62) and (67) with (61), we obtain
@@
$\operatorname{{MSE}}\left(\widehat{F}(Z^{(M)})\right)=\operatorname{bias}\left(\widehat{F}(Z^{(M)})\right)^{2}+\operatorname{Var}\left(\widehat{F}(Z^{(M)}\right),$
@@
$\limsup_{M\to\infty}\dfrac{\operatorname{{MSE}}\left(\widehat{F}(Z^{(M)})\right)}{\frac{a(\sigma)}{|\mathcal{I}_{M}|}}\leq\dfrac{\|F\|_{\infty}K_{1}+K_{2}}{a(\sigma)}+\limsup_{M\to\infty}\dfrac{\operatorname{{MSE}}\left(\widehat{F}(Y^{(M)})\right)}{\frac{a(\sigma)}{|\mathcal{I}_{M}|}}\leq\dfrac{\|F\|_{\infty}K_{1}+K2}{\tau}+1,$

### v3

@@ Appendix A Proofs
The proof consists in proving that
(67) $\limsup_{M\to\infty}|\mathcal{I}_{M}|\left|\operatorname{{MSE}}\left(\widehat{F}(Z^{(M)})\right)-\operatorname{{MSE}}\left(\widehat{F}(Y^{(M)})\right)\right|\leq\|F\|_{\infty}K_{1}+K_{2},$
for some constants $K_{1},K_{2}>0$, which, in view of the convergence rate of $\widehat{F}(Y^{(M)})$ and the assumption $a(\sigma)\geq\tau>0$, implies
$\limsup_{M\to\infty}\dfrac{\operatorname{{MSE}}\left(\widehat{F}(Z^{(M)})\right)}{\frac{a(\sigma)}{|\mathcal{I}_{M}|}}\leq\dfrac{\|F\|_{\infty}K_{1}+K_{2}}{a(\sigma)}+\limsup_{M\to\infty}\dfrac{\operatorname{{MSE}}\left(\widehat{F}(Y^{(M)})\right)}{\frac{a(\sigma)}{|\mathcal{I}_{M}|}}\leq\dfrac{\|F\|_{\infty}K_{1}+K_{2}}{\tau}+1,$
and the conclusion follows.
The proof of (67) is done in four steps. First we use the bias-variance decomposition to reduce the proof of (67) to that of estimating the bias and the variance of $\widehat{F}(Z^{(M)})$. In steps 2 and 3, we estimate the bias and the variance respectively, and in step 4, we assemble everything together to conclude the proof.
Step 1: Bias-variance decomposition. For the i.i.d. measurements ${Y}^{(M)}$, we have the standard bias-variance decomposition
@@
Using the bias-variance decomposition for estimators $\widehat{F}(Z^{(M)})$ and $\widehat{F}(Y^{(M)})$, we obtain
$\left|\operatorname{{MSE}}\left(\widehat{F}(Z^{(M)})\right)-\operatorname{{MSE}}\left(\widehat{F}(Y^{(M)})\right)\right|=$
$\qquad\qquad=\left|\operatorname{bias}(\widehat{F}({Z}^{(M)}))^{2}-\operatorname{bias}(\widehat{F}({Y}^{(M)}))^{2}+\operatorname{Var}(\widehat{F}({Z}^{(M)}))-\operatorname{Var}(\widehat{F}({Y}^{(M)}))\right|$
$\qquad\qquad\leq\left|\operatorname{bias}(\widehat{F}({Z}^{(M)}))^{2}-\operatorname{bias}(\widehat{F}({Y}^{(M)}))^{2}\right|+\left|\operatorname{Var}(\widehat{F}({Z}^{(M)}))-\operatorname{Var}(\widehat{F}({Y}^{(M)}))\right|$
$\qquad\qquad=\operatorname{bias}(\widehat{F}({Z}^{(M)}))^{2}+\left|\operatorname{Var}(\widehat{F}({Z}^{(M)}))-\operatorname{Var}(\widehat{F}({Y}^{(M)}))\right|$
Step 2: Bias estimate. Now, for the dependent data $Z^{(M)}$, we can estimate the bias of the estimator $\widehat{F}(Z^{(M)})$ as
@@
$\leq\underbrace{\left|\dfrac{1}{|\mathcal{I}_{M}|^{2}}\sum_{i\in\mathcal{I}_{M}}\operatorname{Var}\left(F(Z_{i})\right)-\operatorname{Var}\left(\widehat{F}(Y^{(M)})\right)\right|}_{I}$
(70) $\quad+\underbrace{\dfrac{1}{|\mathcal{I}_{M}|^{2}}\sum_{\begin{subarray}{c}(i,j)\in\mathcal{I}_{M}^{2}\\ i\neq j\end{subarray}}\left|\operatorname{Cov}\left(F(Z_{i}),F(Z_{j})\right)\right|}_{II}.$
Step 3.1: Controlling term I. Since ${Y}^{(M)}=\{{Y}_{i}\}_{1\leq i\leq M^{d}}$ are i.i.d., we have
@@
Step 3.2: Controlling term II. Next, we prove that, for any $(i,j)\in\mathcal{I}_{M}^{2}$ with $i\neq j$, it holds that
@@
Step 4: Conclusion. Finally, combining (71) and (76) with (70), we obtain
@@
$\operatorname{{MSE}}\left(\widehat{F}(Z^{(M)})\right)=\operatorname{bias}\left(\widehat{F}(Z^{(M)})\right)^{2}+\operatorname{Var}\left(\widehat{F}(Z^{(M)})\right),$
@@
$\limsup_{M\to\infty}\dfrac{\operatorname{{MSE}}\left(\widehat{F}(Z^{(M)})\right)}{\frac{a(\sigma)}{|\mathcal{I}_{M}|}}\leq\dfrac{\|F\|_{\infty}K_{1}+K_{2}}{a(\sigma)}+\limsup_{M\to\infty}\dfrac{\operatorname{{MSE}}\left(\widehat{F}(Y^{(M)})\right)}{\frac{a(\sigma)}{|\mathcal{I}_{M}|}}\leq\dfrac{\|F\|_{\infty}K_{1}+K_{2}}{\tau}+1,$